# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3.6 Rigorous PP analysis
$\left.\frac{a_{2}\Delta x\Delta y}{\max\left\{6a_{1}\Delta y+2a_{2}\Delta x,6a_{2}\Delta x+2a_{1}\Delta y\right\}\max\left\{\alpha_{2,j+\frac{1}{2}}^{\mu},\alpha_{2,j-\frac{1}{2}}^{\mu}\right\}}\right\},\quad\forall I_{ij}\in\Omega,1\leq\mu\leq q.$
In the above discussions, we restrict ourselves to the first-order forward Euler time discretization. Since an SSP method is a convex combination of the forward Euler method, our PP analysis of the proposed scheme remains valid according to the convexity of $\mathcal{G}$.
@@ 4 Numerical tests
We present some numerical results of proposed schemes for several 1D and 2D MHD problems, to test the accuracy, the ability of capturing shocks, and the robustness in the extreme medium with strong magnetic field and huge discontinuities. The tests below are performed on the uniform 1D meshes and 2D rectangular meshes. We use the HLL flux for all tests. Without loss of generality, we focus on the polynomials of degree $k=2$ for DG solutions with the third-order SSP Runge-Kutta time discretization. The CFL number is set as $0.12$.
@@
This section tests the proposed OEDG scheme through several 1D MHD examples, including a smooth sine wave problem, three shock tube problems, and the MHD version of the Leblanc problem.
@@
We first simulate a smooth MHD problem [44] to validate the accuracy of the PP OEDG scheme in 1D. This problem describes a sine wave propagating with low density, and its exact solution is given by
@@
Periodic boundary conditions are used, and $\gamma=1.4$. Table 1 lists the numerical errors and the corresponding convergence rates at $t=0.1$ for the density. The results show that the 1D $\mathbb{P}^{2}$-based PP OEDG scheme achieves the expected third-order convergence rate.
@@ 4.1.2 Shock tube problems
Next, we simulate three 1D MHD shock tube problems to examine the shock-capturing and non-oscillatory properties of the PP OEDG method. The adiabatic index $\gamma$ is set as $5/3$ for the first two shock tube problems and as $2$ for the third one. We use the third-order PP OEDG scheme with $800$ uniform cells for all the simulations. The initial conditions of the first shock tube problem [56] are given by
@@
and the numerical results at time $t=0.2$ are presented in Figure 2. The second shock tube problem includes a hydrodynamic rarefaction, a switch-on slow shock, a contact discontinuity, a slow shock, a rotational discontinuity, and a fast rarefaction. The initial setups are the same as in [56]:
@@
and Figure 3 shows the numerical results at time $t=0.16$. The third shock tube problem was first suggested by Brio and Wu [57] with the initial data
$\left(\rho,\boldsymbol{u},\boldsymbol{B},p\right)\left(x,0\right)=\left\{\begin{aligned} &\left(1,0,0,0,0.75,1,0,1\right),&x<0,\\ &\left(0.125,0,0,0,0.75,-1,0,0.1\right),&x>0,\\ \end{aligned}\right.$
and the numerical solution at $t=0.1$ is displayed in Figure 4. From these results, one can see that the PP OEDG method correctly captures all the discontinuities such as shocks with high resolution and without any noticeable spurious oscillations.
Figure 2: Numerical solutions of the first MHD shock tube problem at time $t=0.2$ with $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines).
Figure 3: Numerical solutions of the second MHD shock tube problem at time $t=0.16$ with $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines).
Figure 4: Numerical solutions of the third MHD shock tube problem at time $t=0.1$ computed by $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines).
@@ 4.1.3 Leblanc problem
In order to test the robustness of the 1D PP OEDG scheme, we run a challenging MHD Riemann problem with a strong magnetic field, a large jump of the initial pressure, and extremely low plasma-beta ($\beta=4\times 10^{-8}$). This problem is a variant of the Leblanc problem [37] of the gas dynamics, and was proposed in [44]. The initial setting is as follows:
@@
The computational domain is set to $\left[-10,10\right]$, and the adiabatic index $\gamma$ is taken as $1.4$. Figure 5 displays the numerical results of density and magnetic pressure at $t=0.00003$. As expected, negative density or pressure does not appear in the cell averages of solution, indicating that the 1D PP OEDG scheme exhibits good stability.
Figure 5: Density and magnetic pressure of the 1D Leblanc problem at time $t=0.00003$ with $2000$ cells (symbols $\circ$”) and $10000$ cells (solid lines).
@@
This section conducts six 2D MHD tests using $\mathbb{P}^{2}$–based PP LDF OEDG method, including a smooth sine wave problem, the Orszag–Tang problem, rotor problem, shock cloud interaction, MHD blast wave, and MHD jet flows.
@@
To verify the convergence order of the PP LDF OEDG method, we simulate a 2D smooth sine wave problem [43], which describes a MHD sine wave propagating with low density. The exact solution of this problem is given by
@@
and the adiabatic index $\gamma=1.4$. Table 2 provides the numerical errors and corresponding convergence rates of density at time $t=0.1$. The results show that the $\mathbb{P}^{2}$–based PP LDF OEDG scheme achieves the third-order accuracy overall.
@@ 4.2.2 Orszag–Tang problem
We now consider a widely-used Orszag–Tang vortex problem [58]. For this problem, the vortex starts from a smooth initial field configuration and quickly forms shocks before transitioning to 2D MHD turbulence. The initial conditions are as follows:
@@
where the adiabatic index $\gamma$ is taken as $5/3$. The computational domain is set as $\left[0,2\pi\right]\times\left[0,2\pi\right]$, and the periodic boundary conditions are used on all boundaries.
Figure 6 shows the density and thermal pressure obtained by third-order PP LDF OEDG scheme at $400\times 400$ resolution. We can clearly see both sharp shocks and smooth flows, which are in good agreement with those in [59, 60]. For the Orszag–Tang vortex problem, Jiang and Wu [61] previously reported that the negative pressure was occurred at $t\approx 3.9$ in their test. In our implementation for third-order PP LDF OEDG scheme, negative pressure was not observed. To further verify the convergence of the scheme, Figure 7 plots the profiles of density and thermal pressure along the line of $y=0.625\pi$ at time $t=3$ under $200\times 200$ and $400\times 400$ resolutions. In Figure 7, the shock discontinuities for density and thermal pressure are formed near $x=0.4,0.5,1.6$ and $4.4$, which is consistent with the results in [60, 26]. According to the comparison of second-order and fourth-order schemes for pressure profiles by Yang et al. [60], the high accuracy scheme results in sharper shock discontinuities at $x=0.4$ and $x=0.5$. In Figure 7, it is observed that the shock discontinuities around $x=0.4$ and $x=0.5$ with $400\times 400$ resolution are much sharper than those with $200\times 200$ resolution, verifying the convergence of our scheme.
Figure 6: Maps of (a): density and (b): thermal pressure for Orszag-Tang vortex problem at time $t=3$ with $400\times 400$ cells.
Figure 7: Profiles of density (left) and thermal pressure (right) along line $y=0.625\pi$ for Orszag-Tang vortex problem at time $t=3$ with $200\times 200$ cells (red lines) and $400\times 400$ cells (black lines).
@@ 4.2.3 Rotor problem
We next consider the rotor problem [21] to further test the performance of $\mathbb{P}^{2}$–based PP LDF OEDG method. In this rotor problem, a dense disk of fluid is placed in a static fluid background, with a gradual velocity tapering layer between the dense disk edge and the static ambient fluid, and the disk will rotate within ambient fluid over time. The initial setup is given by
@@
where $r_{0}=0.1$, $r_{1}=0.115$, $r=\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}$, and $f=(r_{1}-r)/(r_{1}-r_{0})$. The adiabatic index $\gamma$ is taken as $5/3$. The computational domain is $\left[0,1\right]\times\left[0,1\right]$, and periodic boundary conditions are used.
@@
In Figure 8, the contours of density, thermal pressure, Mach number and magnetic pressure are presented at $t=0.295$ for the third-order PP LDF OEDG method with $400\times 400$ cells. We can see that the central contours show good conservation of the circular rotation pattern, and such performance was found by Tóth [62] to be challenging for some other MHD numerical schemes. According to Li and Shu [7], the numerical solution around central part of the Mach number existed some distortion, which was attributed to the divergence error in the magnetic field. Figure 9 zooms in the central part of the Mach number at different resolutions. No distortions are observed at the outskirts of the central almond-shaped disk region, which suggests that the divergence error in the magnetic field is well controlled by our LDF OEDG scheme. To further examine the performance of our scheme, Figure 10 presents two slices of the magnetic field at time $t=0.295$ across horizontal and vertical cuts through the centre of the box with different mesh resolutions. The results show good agreement with previous studies [26, 63]. It was found that in [26] higher-order scheme exhibits higher peaks near $x=0.35,0.65$ of $B_{1}$ and $x=0.325,0.675$ of $B_{2}$, as well as sharper shock discontinuities around $x=0.4,0.6$ of $B_{1}$. In Figure 10, the same trend is shown in the high-resolution results compared to the low-resolution results, which reflects the convergence of our scheme. According to Liu et al. [26], the overshoots at $y=0.05,0.95$ for $B_{1}$ and at $x=0.38,0.62$ for $B_{2}$ were generated by their third-order finite volume scheme. Figure 10 shows that no overshoot is observed, verifying that the PP LDF OEDG scheme can suppress oscillations well.
@@
Figure 8: Contour maps of (a) density $\rho$, (b) thermal pressure $p$, (c) Mach number $|\boldsymbol{u}|/c$, and (d) magnetic pressure $|\boldsymbol{B}|^{2}/2$ at time $t=0.295$ with $400\times 400$ cells for rotor problem. Twenty-five equally spaced contours are used for density with range $[0.8,10.2]$, thermal pressure with range $[0.01,1.1]$, Mach number with range $[0,2.2]$, and magnetic pressure with range $[0.02,0.68]$.
@@
Figure 9: Zoom-in central part of Mach number at time $t=0.295$ for rotor problem. The left panel is from $200\times 200$ cells, and the right panel is from $400\times 400$ cells. Forty equally spaced contours are utilized within the range $[0,2.2]$.
@@
Figure 10: Slices of $B_{1}$ (left) and $B_{2}$ (right) along the lines $x=0.5$ and $y=0.5$ for rotor problem at time $t=0.295$ with $200\times 200$ cells (red lines) and $400\times 400$ cells (black lines).
4.2.4 Shock cloud interaction
@@
To further validate the shock-capturing and PP performance of our scheme, we consider shock cloud interaction [64]. This example involves strong MHD shocks interacting with a disruption of dense cloud. The initial conditions are given by the left and right sates separated by a discontinuity along the line $x=0.6$:
@@
and a circular cloud with a radius of $0.15$ centered on $(0.8,0.5)$ is placed on the right side of the discontinuity. The cloud has the same setups as the ambient fluid except for a higher density $10$. Our computational domain is $[0,1]\times[0,1]$, and adiabatic index $\gamma$ is $5/3$. For this problem, supersonic inflow condition is employed on the right boundary, and outflow conditions are utilized on the others.
Figure 11: Maps of (a) density $\rho$, (b) thermal pressure $p$, (c) velocity $|\boldsymbol{u}|$, and (d) magnetic pressure $|\boldsymbol{B}|^{2}/2$ at time $t=0.06$ with $400\times 400$ cells for shock cloud interaction.
Figure 11 displays numerical results of density, thermal pressure, velocity, and magnetic pressure at time $t=0.06$ with $400\times 400$ cells for our third-order PP LDF OEDG scheme. One can see that the complex flow structures and discontinuities are well captured with high resolution, and the results show good agreement with those in the literature, e.g., [62, 65, 43, 44]. We found that once the OE procedure is turned off in this test, the code will break down and cause simulation failure, which proves the stability of our scheme. In addition, we also observed that negative pressure will appear in the cell average of numerical solutions if we remove either the Godunov–Powell source term or the PP limiter. This further confirms the importance of the discretized Godunov–Powell source term and the PP limiter for preserving positivity.

### v2

@@ 3.6 Rigorous PP analysis
$\bar{\omega}_{\star}(\delta,\mathbb{P}^{k})=\left[\frac{14}{3}+\frac{2}{3}\sqrt{78\,\delta^{2}+46}\cos\left(\frac{1}{3}\arccos\frac{1476\,\delta^{2}-244}{(78\,\delta^{2}+46)^{\frac{3}{2}}}\right)\right]^{-1}.$
For higher degree $k$; see [58] for more details.
Thanks to the “two-state” inequalities established in Lemma 3 and the optimal convex decomposition, our above PP CFL condition (2) is notably improved, compared to the estimates in [48] for general meshes when reduced to Cartesian meshes.
In our above analysis, our focus has been on employing the forward Euler time discretization. Given that a high-order SSP method is essentially a convex combination of the forward Euler approach, our PP analysis of the proposed schemes persists as applicable for high-order SSP time discretization owing to the convexity $\mathcal{G}$.
@@ 4 Numerical tests
This section gives several 1D and 2D MHD numerical examples to validate the accuracy, the essentially non-oscillatory shock-capturing capability, and the robustness of the proposed PP LDF OEDG method on uniform 1D meshes and 2D rectangular meshes. Without loss of generality, we focus on the third-order ($k=2$) OEDG method, coupled with the third-order SSP Runge–Kutta time discretization. We use the HLL flux and set the CFL number to $0.12$.
@@
This subsection assesses the OEDG scheme through various 1D MHD examples, including a smooth sine wave problem, three shock tube problems, and the MHD version of the Leblanc problem.
@@
We begin by simulating a smooth MHD sine wave problem to evaluate the accuracy of the PP OEDG scheme in the 1D setting. This problem involves a sine wave propagating with low density, and its exact solution is described by:
@@
Periodic boundary conditions are applied, and $\gamma=1.4$. Table 1 lists the numerical errors and the corresponding convergence rates at $t=0.1$ for the density. The results confirm that the 1D $\mathbb{P}^{2}$-based PP OEDG scheme achieves the expected third-order convergence rate. This confirms that the OE procedure and PP limiter do not compromise the optimal convergence order of the DG solutions.
@@ 4.1.2 Shock tube problems
To evaluate the shock-capturing and non-oscillatory properties of the PP OEDG method, we simulate three 1D MHD shock tube problems. For the first two problems, the adiabatic index $\gamma$ is set to $5/3$, and for the third problem, it is set to $2$. All simulations employ the third-order PP OEDG scheme with $800$ uniform cells. The initial conditions of the first shock tube problem [59] are given by
@@
and the numerical results at time $t=0.2$ are presented in Figure 3. The second shock tube problem encompasses a hydrodynamic rarefaction, a switch-on slow shock, a contact discontinuity, a slow shock, a rotational discontinuity, and a fast rarefaction, following the setup in [59]:
@@
and Figure 4 shows the numerical results at time $t=0.16$. The third shock tube problem, initially proposed by Brio and Wu [60], features the following initial conditions:
$\left(\rho,\boldsymbol{u},\boldsymbol{B},p\right)\left(x,0\right)=\left\{\begin{aligned} &\left(1,0,0,0,0.75,1,0,1\right),&x<0,\\ &\left(0.125,0,0,0,0.75,-1,0,0.1\right),&x>0.\\ \end{aligned}\right.$
The numerical results at time $t=0.1$ are shown in Figure 5. From these results, one can see that the PP OEDG method accurately captures all the discontinuities such as shocks with high resolution and without any noticeable spurious oscillations.
Figure 3: Numerical solutions of the first MHD shock tube problem at time $t=0.2$ with $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines). Left column from top to bottom: $\rho$, $u_{1}$, $u_{3}$, $B_{3}$. Right column from top to bottom: $p$, $u_{2}$, $B_{2}$, $E$.
Figure 4: Numerical solutions of the second MHD shock tube problem at time $t=0.16$ with $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines). Left column from top to bottom: $\rho$, $u_{1}$, $u_{3}$, $B_{3}$. Right column from top to bottom: $p$, $u_{2}$, $B_{2}$, $E$.
Figure 5: Numerical solutions of the third MHD shock tube problem at time $t=0.1$ computed by the third-order OEDG scheme with $800$ cells (symbols $\circ$”) and $4000$ cells (solid lines). Left column from top to bottom: $\rho$, $u_{1}$, $B_{2}$. Right column from top to bottom: $p$, $u_{2}$, $E$.
@@ 4.1.3 Leblanc problem
To assess the robustness of the 1D PP OEDG scheme, we test a challenging MHD Riemann problem with a strong magnetic field, a large pressure jump, and an extremely low plasma-beta ($\beta=4\times 10^{-8}$). This problem is a variant of the Leblanc problem [40] of the gas dynamics and was proposed in [48]. The initial conditions are defined as follows:
@@
The computational domain spans $\left[-10,10\right]$, and the adiabatic index $\gamma$ is set to $1.4$. Figure 6 presents the numerical results for density and magnetic pressure at $t=0.00003$. As anticipated, no negative density or pressure values are observed in the OEDG solution, demonstrating the robustness and stability of the 1D PP OEDG scheme.
Figure 6: Density logarithm (left) and magnetic pressure (right) of the 1D Leblanc problem at time $t=0.00003$ with $2000$ cells (symbols $\circ$”) and $10000$ cells (solid lines).
@@
This subsection conducts six 2D MHD tests on the third-order PP LDF OEDG method, which include simulations of a smooth sine wave, the Orszag–Tang vortex, a rotor problem, shock-cloud interaction, an MHD blast wave, and MHD jet flows.
@@
To evaluate the convergence order of the PP LDF OEDG method, we simulate a 2D smooth sine wave problem [47], which involves an MHD sine wave propagating through a medium with low density. The exact solution of this problem is expressed as follows:
@@
and the adiabatic index $\gamma=1.4$. Table 2 presents the numerical errors and corresponding convergence rates for density at time $t=0.1$. The results demonstrate that the $\mathbb{P}^{2}$-based PP LDF OEDG scheme consistently achieves third-order accuracy. That is, the optimal convergence rate is not destroyed by the LDF OE procedure and the PP limiter.
@@ 4.2.2 Orszag–Tang problem
We next examine the benchmark Orszag–Tang vortex problem [61]. This problem begins with smooth initial data, rapidly forming shocks, and eventually transitioning into 2D MHD turbulence. The initial conditions are set as
@@
where the adiabatic index $\gamma$ is $5/3$. The computational domain spans $\left[0,2\pi\right]\times\left[0,2\pi\right]$, with periodic boundary conditions applied on all edges.
Figure 7 displays contour plots of density, magnetic pressure, and Mach number at times $t=3$ and $t=4$, obtained by the third-order PP LDF OEDG scheme with $400\times 400$ cells. Our method accurately resolves the shocks and smooth flows, which are in good agreement with those in [7, 62, 10]. For this problem, Jiang and Wu [63] previously reported that negative pressure was occurred at $t\approx 3.9$ in their computation. Thanks to the PP property of our scheme, negative pressure was not observed throughout our simulation. To further verify the convergence of our scheme, Figure 8 plots the profiles of density and thermal pressure along the line of $y=0.625\pi$ at time $t=3$ under $200\times 200$ and $400\times 400$ resolutions. As shown in Figure 8, the shock discontinuities for density and thermal pressure are formed near $x=0.4,0.5,1.6$ and $4.4$, which are consistent with the results in [64, 28]. As discussed by Yang et al. in [64], high-resolution simulations typically exhibit sharper shock profiles, particularly noticeable at $x=0.4$ and $x=0.5$. In Figure 8, it is observed that the shock discontinuities around $x=0.4$ and $x=0.5$ with $400\times 400$ resolution are much sharper than those with $200\times 200$ resolution, indicating the convergence and effectiveness of our scheme.
Figure 7: Contour plots for the Orszag–Tang problem with $400\times 400$ cells. From top to bottom: density $\rho$, magnetic pressure $\|\boldsymbol{B}\|^{2}/2$, and Mach number $\|\boldsymbol{u}\|/c$. Left: $t=3$; right: $t=4$.
Figure 8: Profiles of density (left) and thermal pressure (right) along line $y=0.625\pi$ at time $t=3$ for the Orszag–Tang problem with $200\times 200$ cells (dash dot lines) and $400\times 400$ cells (solid lines).
@@ 4.2.3 Rotor problem
We further evaluate the PP LDF OEDG method by simulating the rotor problem [22]. This problem involves a dense disk of fluid embedded in a static fluid background, with a velocity tapering layer between the dense disk’s edge and the static ambient fluid. As time progresses, the disk rotates within the ambient fluid. The initial configurations are set as follows:
@@
where $r_{0}=0.1$, $r_{1}=0.115$, $r=\sqrt{(x-0.5)^{2}+(y-0.5)^{2}}$, and $f=(r_{1}-r)/(r_{1}-r_{0})$. The adiabatic index $\gamma$ is $5/3$, and the computational domain is $\left[0,1\right]\times\left[0,1\right]$ with periodic boundary conditions.
@@
Figure 9 presents the contours of density, thermal pressure, Mach number, and magnetic pressure at $t=0.295$, computed using the third-order PP LDF OEDG method with $400\times 400$ cells. The contours show good conservation of the circular rotation pattern, which Tóth [65] found challenging for some other MHD numerical schemes. Distortions may appear in the numerical solution around the central part of the Mach number if there are large divergence errors in the magnetic field, as observed in [6]. However, our simulation results show no distortions around the central almond-shaped disk region, suggesting effective control of divergence errors by our LDF OEDG scheme. To further assess the performance of our scheme, Figure 10 zooms into the central part of our computed Mach number at different resolutions, showing no observed distortions. Figure 11 presents two slices of the magnetic field at time $t=0.295$ across horizontal and vertical cuts through the center of the domain at different mesh resolutions, with results that align well with those reported by [28, 66]. It was found that in [28] higher-order scheme exhibits higher peaks near $x=0.35,0.65$ of $B_{1}$ and $x=0.325,0.675$ of $B_{2}$, as well as sharper shock discontinuities around $x=0.4,0.6$ of $B_{1}$. In Figure 11, the same trend is observed in our high-resolution results compared to the low-resolution ones, reflecting the convergence behavior of our scheme. According to Liu et al. [28], the overshoots at $y=0.05,0.95$ for $B_{1}$ and at $x=0.38,0.62$ for $B_{2}$ were generated by their third-order finite volume scheme. Our results in Figure 11 show no overshoots, further verifying that the LDF OE procedure effectively suppresses nonphysical oscillations.
@@
Figure 9: Contour plots of density $\rho$ (top left), thermal pressure $p$ (top right), Mach number $\|\boldsymbol{u}\|/c$ (bottom left) and magnetic pressure $\|\boldsymbol{B}\|^{2}/2$ (bottom right) at time $t=0.295$ for the rotor problem with $400\times 400$ cells. Twenty-five equally spaced contours are used for density with range $[0.8,10.2]$, thermal pressure with range $[0.01,1.1]$, Mach number with range $[0,2.2]$, and magnetic pressure with range $[0.02,0.68]$.
@@
Figure 10: Zoom-in central parts of Mach number at time $t=0.295$ for the rotor problem with $200\times 200$ cells (left) and $400\times 400$ cells (right). Forty equally spaced contours are utilized within the range $[0,2.2]$.
@@
Figure 11: Slices of $B_{1}$ (left) and $B_{2}$ (right) along the lines $x=0.5$ and $y=0.5$ at time $t=0.295$ for the rotor problem with $200\times 200$ cells (dash dot lines) and $400\times 400$ cells (solid lines).
4.2.4 Shock-cloud interaction
@@
To further validate the shock-capturing and PP capabilities of our structure-preserving OEDG scheme, we simulate the shock-cloud interaction problem [67]. This test involves strong MHD shocks interacting with a dense cloud, which is disrupted by the interaction. The initial conditions feature left and right states separated by a discontinuity along the line $x=0.6$:
@@
and includes a circular cloud with a radius of $0.15$ centered at $(0.8,0.5)$ on the right side of the discontinuity. The cloud has the same state as the surrounding medium but with a higher density of $10$. The computational domain is $[0,1]\times[0,1]$, and adiabatic index $\gamma$ is $5/3$. For this problem, supersonic inflow condition is employed on the right boundary, while outflow conditions are utilized on the others.
Figure 12: The density $\rho$ (top left), thermal pressure $p$ (top right), velocity $\|\boldsymbol{u}\|$ (bottom left) and magnetic pressure $\|\boldsymbol{B}\|^{2}/2$ (bottom right) at time $t=0.06$ for the shock cloud interaction with $400\times 400$ cells.
Figure 12 displays the numerical results at time $t=0.06$ for density, thermal pressure, velocity magnitude, and magnetic pressure, computed using our third-order PP LDF OEDG scheme on a $400\times 400$ mesh. The results successfully capture complex flow structures and discontinuities at high resolution, demonstrating good agreement with those in the literature, e.g., [65, 68, 47, 48]. We note that disabling the LDF OE procedure in this simulation leads to failure, underscoring the critical role of our OE technique. Furthermore, removal of either the Godunov–Powell source terms or the PP limiter resulted in negative pressure values in the cell averages of numerical solutions, reaffirming the necessity of these elements for ensuring positivity in the simulation.